OpenStudy (anonymous):

Find the limit, if it exists. (If an answer does not exist, enter DNE.) lim sqroot t+ t^2/6t − t^2 t→inf

OpenStudy (turingtest):

\[\lim_{t\to\infty}\frac{\sqrt{t+t^2}}{6t-t^2}\]is this right?

OpenStudy (anonymous):

the sqaure root is on for the t

OpenStudy (turingtest):

\[\lim_{t\to\infty}\frac{\sqrt t+t^2}{6t-t^2}\]

OpenStudy (anonymous):

yeyeyeey

OpenStudy (turingtest):

divide top and bottom by t^2

OpenStudy (anonymous):

\[\frac{ \sqrt{t} }{ 6t }\]

OpenStudy (turingtest):

?

OpenStudy (turingtest):

divide each term by t^2

OpenStudy (anonymous):

i'm oo confused can you show me

OpenStudy (turingtest):

\[\lim_{t\to\infty}\frac{\frac1{t^2}(\sqrt t+t^2)}{\frac1{t^2}(6t-t^2)}\]

OpenStudy (turingtest):

distribute

OpenStudy (anonymous):

whats is the answer

OpenStudy (turingtest):

this site is about teaching, not handing out answers can you not simplify the numerator and denominator? this is algebra, you should know this.

OpenStudy (anonymous):

sqroot(t)/t^2

OpenStudy (anonymous):

6t;t^2

OpenStudy (turingtest):

what about the t^2's on the right?

OpenStudy (turingtest):

what is\[\frac1{t^2}(\sqrt t+t^2)\]?

OpenStudy (anonymous):

doesn't the t^2/t^2 cancels out

OpenStudy (turingtest):

cancels to what?

OpenStudy (anonymous):

1

OpenStudy (turingtest):

yes, but you never wrote the 1

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

(sqroot (t)+1)/t^2/(6t-1)/t^2

OpenStudy (turingtest):

those 1's should not be over t^2, they cancel like you said

OpenStudy (turingtest):

just do the top\[\frac1{t^2}(\sqrt t+t^2)\]

OpenStudy (anonymous):

okaay

OpenStudy (anonymous):

sqroot (t)/t^2 +1/ 6t/t^2 +1

OpenStudy (turingtest):

yes, and what is sqrt(t)/t^2 simplified?

OpenStudy (anonymous):

i don"t know

OpenStudy (turingtest):

\[\sqrt[b]{x^a}=x^{a/b}\]so\[\sqrt t=t^{1/2}\]and\[\frac{t^a}{t^b}=t^{a-b}\]use those two rules together

OpenStudy (anonymous):

okaay thambi

OpenStudy (anonymous):

3root (t^2)